John hosts an art workshop on the weekends. He has an average of 14 students in each session and charges a fee of $12 per session. He estimates that for every $2 increases in the fee, the average number of students reduces by 1.
Complete the equation that models this scenario, where c(x) is the revenue generated and x is the number of $2 fee increases.

Answer by stanbon(75887)

John hosts an art workshop on the weekends. He has an average of 14 students in each session and charges a fee of $12 per session. He estimates that for every $2 increases in the fee, the average number of students reduces by 1.
------
Original Revenue DATA:
fee = 12
# of students = 14
Revenue = 12*14 = $168
---------
Hypothesized Revenue ::
fee = 12 + 2x
# of students = 14-x
---
Ans:: C(x) = (12+2x)(14-x)
Complete the equation that models this scenario, where c(x) is the revenue generated and x is the number of $2 fee increases.
--------------
Cheers,
Stan H.
-----------